# Entropy formula for surface diffeomorphisms

### v1

@@ 4.2 Decomposition of the volume growth
We claim that $\mu$ must have a unique positive Lyapunov exponent (hence $\hat{\mu}$ is the unique unstable lift of $\mu$), due to the choice of $m_{n}$. Indeed, if not, $\mu$ would have an ergodic component with two positive Lyapunov exponents and this ergodic component must support on some periodic source $z$ with $\mu$ positive measure ( see [21, Proposition 4.4] or [8, Fact in Claim 7.3]). Let $r>0$ be small enough such that if $y\notin B\left(f^{i}(z),\frac{\varepsilon_{f}}{100\left\|Df\right\|}\right)$ for some $i\in\mathbb{N}$, then $f^{k}(y)\notin B(z,r)$ for all $k\in\mathbb{N}$  [^21] This is possible since $z$ is a periodic source. Let $T_{z}$ denote the period of $z$. We can find a small number $0<\delta<\frac{\varepsilon_{f}}{100\left\|Df\right\|}$ such that for $0\leq k\leq T_{z}-1$, $f^{-T_{z}}\left(B\left(f^{k}(z),\delta\right)\right)\subset B\left(f^{k}(z),\delta\right).$ Then choose $r>0$ small enough such that for $0\leq k\leq T_{z}-1$, $f^{k}\left(B\left(z,r\right)\right)\subset B\left(f^{k}(z),\delta\right).$ . Since $\mu$ is not zero measure, we may assume for all large $n$,
$m_{n}<n.$
Since $x_{n,p}\in\sigma_{n}^{m_{n}}$, by the definition of $\sigma_{n}^{m_{n}}$, there must exist some $0\leq i<m_{n}+1$ such that
$f^{i}(x_{n,p})\notin B\left(f^{i}(z),\frac{\varepsilon_{f}}{100\left\|Df\right\|}\right).$
@@
which is a contradiction. Hence $\mu$ must have a unique positive Lyapunov exponent and $\hat{\mu}$ is the unique unstable lift of $\mu$.
@@
Now we can use Proposition 4.9 to prove Theorem 2.1.
@@
Define
$\Phi(\alpha):=\alpha\cdot h_{\rm top}(f)+(1-\alpha)\cdot\frac{\log\left\|Df\right\|}{r}+4\log\left(\log\left\|Df\right\|+\log\left\|Df^{-1}\right\|+2\right)+C_{r},\qquad\alpha\in[0,1]$
where $C^{r}$ is the universal constant depending only $r$ from Proposition 4.9.
@@
$\limsup_{n\to+\infty}\frac{1}{n}\log\sup_{\sigma}{\rm Vol}\left(f^{n}(\sigma)\right)\leq\max_{\alpha\in[0,1]}\Phi(\alpha).$ (19)
@@
$\limsup_{n\to+\infty}\frac{1}{n}\log\sup_{\sigma}{\rm Vol}\left(f^{n}(\sigma)\right)>\max_{\alpha\in[0,1]}\Phi(\alpha).$
@@
$\limsup_{k\to+\infty}\frac{1}{n_{k}}\log{\rm Vol}\left(f^{n_{k}}(\sigma_{n_{k}})\right)>\max_{\alpha\in[0,1]}\Phi(\alpha).$
Since each $\sigma_{n_{k}}$ can be decomposed into a uniformly bounded number of $\varepsilon_{f}$-bounded curves (see Lemma 6.3), we may, without loss of generality, assume that the sequence $\{\sigma_{n_{k}}\}$ itself consists of $\varepsilon_{f}$-bounded curves. Hence, by Proposition 4.9, there exists $\alpha\in[0,1]$ such that
$\limsup_{k\to+\infty}\frac{1}{n_{k}}\log{\rm Vol}\left(f^{n_{k}}(\sigma_{n_{k}})\right)\leq\Phi(\alpha)\leq\max_{\alpha\in[0,1]}\Phi(\alpha),$
@@
$\limsup_{n\to+\infty}\frac{1}{n}\log\sup_{\sigma}{\rm Vol}\left(f^{n}(\sigma)\right)\leq\max_{\alpha\in[0,1]}\left\{\alpha\cdot h_{\rm top}(f)+(1-\alpha)\cdot\frac{\lambda^{+}(f)}{r}\right\}.$
@@ 5 Volume growth of the tangent cocycle
In this section, we study the relationship between the volume growth of the tangent cocycle and entropy. More precisely, we establish the chain of inequalities
$h_{\rm top}(f)\leq\liminf_{n\to+\infty}\frac{1}{n}\log\int_{M}\left\|Df^{n}_{x}\right\|\,dx\leq\limsup_{n\to+\infty}\frac{1}{n}\log\int_{M}\left\|Df^{n}_{x}\right\|\,dx\leq\alpha\cdot h_{\rm top}(f)+(1-\alpha)\cdot\frac{\lambda^{+}(f)}{r}.$
@@ 5.1 Upper bound of the volume growth rate
The following lemma asserts the existence of a point at which the exponential growth rate of the integral of $\left\|Df^{n}_{x}\right\|$ is fully realized locally. In other words, the global asymptotic growth of $\int_{M}\left\|Df^{n}_{x}\right\|\,dx$ can be detected on arbitrarily small neighborhoods of a single point.
@@
Let
$\lambda(S):=\limsup_{n\to+\infty}\frac{1}{n}\log\int_{S}\left\|Df^{n}_{x}\right\|\,dx.$
@@
$\limsup_{n\to\infty}\frac{1}{n}\log\left(\sum_{i=1}^{k}a_{n,i}\right)=\max_{1\leq i\leq k}\left(\limsup_{n\to\infty}\frac{1}{n}\log a_{n,i}\right),$
@@
$\lambda(M)=\lambda\left(\bigcup_{i=1}^{k}U_{z_{i}}\right)=\max_{1\leq i\leq k}\lambda(U_{z_{i}}).$
Since $\lambda(U_{z_{i}})<\lambda(M)$ for all $i$, we obtain $\lambda(M)<\lambda(M)$, a contradiction. ∎
The main idea to prove Theorem 2.2 is to reduce the estimation of the volume growth of the tangent cocycle to the estimation of the volume growth of the one-dimensional sub-manifolds. Using Fubini’s theorem, the two-dimensional integral of $\left\|Df_{x}^{n}\right\|$ is decomposed into integrals along horizontal and vertical lines, reducing the problem to estimating the volume growth of one-dimensional curves. The conclusion then follows by applying the curve volume growth estimate from Proposition 4.9.
@@
Let $\mathcal{A}:=\left\{\left(U_{i},\varphi_{i}\right):~\varphi_{i}:U_{i}\to[0,1]^{2}\right\}_{0\leq i<N_{0}}$ be a finite family of charts that cover $M$. Let $z$ be the point in Lemma 5.1. Without loss of generality, we assume $z\in U_{0}$ and $U_{0}$ is a ball at $z$ with diameter far less than $\varepsilon_{f}$ ($\varepsilon_{f}$ is from Proposition 4.1) such that for any line $\sigma$ in $0,1]^{2}$, $\varphi^{-1}_{0}\left(\sigma\right)$ is an $\varepsilon_{f}$-bounded curve on $M$. Let $K$ be a sufficiently large number (independent of $n$) such that
  • for any $0\leq i<N_{0}$ and any curve $\sigma$ on $U_{i}$,
@@
  • for any $i$ and $j$, assuming $x\in U_{i}$ and $f^{n}(x)\in U_{j}$,
    $K^{-1}\cdot\left\|D\left(\varphi_{j}\circ f^{n}\circ\varphi^{-1}_{i}\right)_{\varphi_{i}(x)}\right\|\leq\left\|Df_{x}^{n}\right\|\leq K\cdot\left\|D\left(\varphi_{j}\circ f^{n}\circ\varphi^{-1}_{i}\right)_{\varphi_{i}(x)}\right\|.$
Given $n\in\mathbb{N}$, we fix any $0\leq i<N_{0}$ such that
$\int_{U_{0}^{i}}\left\|Df^{n}_{x}\right\|\,dx\geq\frac{1}{N_{0}}\int_{U_{0}}\left\|Df^{n}_{x}\right\|\,dx$ (20)
where
$U_{0}^{i}:=\left\{x\in U_{0}:~f^{n}\left(x\right)\in U_{i}\right\}.$
Let $\left(x_{1},x_{2}\right)$ be the standard coordinates in $\mathbb{R}^{2}$. By the choice of $K$, we note that
@@
where
$F^{n}:=\varphi_{i}\circ f^{n}\circ\varphi^{-1}_{0}.$
@@
Applying Proposition 4.9 to $\sigma_{n}$, recalling (20) and (21), we have
$\limsup_{n\to+\infty}\frac{1}{n}\log\int_{M}\left\|Df^{n}_{x}\right\|\,dx$ $=\limsup_{n\to+\infty}\frac{1}{n}\log\int_{U_{0}}\left\|Df^{n}_{x}\right\|\,dx$ (22)
@@
$\leq\max_{\alpha\in[0,1]}\left\{\alpha\cdot h_{\rm top}(f)+\left(1-\alpha\right)\cdot\frac{\log\left\|Df\right\|}{r}\right\}$
$\quad+4\log\left(\log\left\|Df\right\|+\log\left\|Df^{-1}\right\|+2\right)+C_{r}$
where $C_{r}$ is a universal constant depending only on $r$.
Note that
$\limsup_{n\to+\infty}\frac{1}{n}\log\int_{M}\left\|Df^{n}_{x}\right\|\,dx=\limsup_{n\to+\infty}\frac{1}{qn}\log\int_{M}\left\|Df^{qn}_{x}\right\|\,dx,\quad h_{\rm top}(f)=\frac{h_{\rm top}(f^{q})}{q}.$
Replacing $f$ by $f^{q}$ and putting $\frac{1}{q}$ on both sides in (22), by the arbitrariness of $q$, we get the conclusion:
$\limsup_{n\to+\infty}\frac{1}{n}\log\int_{M}\left\|Df^{n}_{x}\right\|\,dx\leq\max_{\alpha\in[0,1]}\left\{\alpha\cdot h_{\rm top}(f)+(1-\alpha)\cdot\frac{\lambda^{+}(f)}{r}\right\}.$
@@ 5.2 Lower bound of the volume growth rate
and whose cardinality is minimal among all finite subsets of $M$ with this property. Recall (see [25] for basic background in ergodic theory) that a subset $S\subset M$ is called a $(n,\varepsilon)$ separated set if for any $x,y\in S$ with $x\neq y$, there is $0\leq k<n$ such that $d(f^{k}(x),f^{k}(y))>\varepsilon$. Moreover, a separated set $S$ is called maximal if its cardinality is larger than or equal to the cardinality of any other separated set. Given a subset $\Omega$, let $S(n,\varepsilon,\Omega)$ denote a maximal $(n,\varepsilon)$ separated set of $\Omega$. By maximality, the union of the dynamical balls $\{B(x,n,\varepsilon)\}_{x\in S(n,\varepsilon,\Omega)}$ covers $\Omega$.
@@
The following lemma, due to Przytycki, as an application of Pesin theory, provides a uniform lower bound on the integral of the derivative over dynamical balls centered at typical points of a hyperbolic measure.
@@
[22, Lemma 2.1] Let $f$ be a $C^{r}$ ($r>1$) diffeomorphism on a compact surface $M$ and let $\mu$ be a hyperbolic ergodic measure. There are a subset $\Omega\subset M$ with $\mu\left(\Omega\right)>\frac{1}{2}$ and a constant $C>0$ such that for any $z\in\Omega$, any $n\in\mathbb{N}$ and any $0<\varepsilon<2026$,
@@
$\int_{M}\left\|Df_{x}\right\|\,dx$ $\geq\#S(n,2\varepsilon,\Omega)\cdot\min_{z\in S(n,2\varepsilon,\Omega)}\int_{B(z,n,\varepsilon)}\left\|Df_{x}\right\|\,dx$
@@
$\liminf_{n\to+\infty}\frac{1}{n}\log\int_{M}\left\|Df_{x}\right\|\,dx\geq\liminf_{n\to+\infty}\frac{1}{n}\log\#S(n,2\varepsilon,\Omega)-\log 2.$
Since $\mu(\Omega)>\frac{1}{2}$, we have $S(n,2\varepsilon,\Omega)\geq\#S_{\frac{1}{2}}(n,2\varepsilon)$. By Lemma 5.3, by the arbitrariness of $\varepsilon$, we then have
@@ 6 Appendix
The following lemma establishes an upper semi-continuity property for integrals of sub-additive potentials along converging sequences of invariant measures. This result is a special case of [11, Lemma 2.3].
@@
Let $X$ be a compact metric space and let $\mu$ be a Borel probability measure on $X$. Let $\{f_{n}\}_{n\geq 1}$ be a sequence of measurable functions on $X$ such that $\sup_{n\geq 1}\left\|f_{n}\right\|_{\infty}<+\infty$. Then
$\int_{X}\limsup_{n\to+\infty}f_{n}\,d\mu\geq\limsup_{n\to+\infty}\int_{X}f_{n}\,d\mu.$
@@
Take $N=\left\lceil 1/\varepsilon\right\rceil$ and define the step size $\delta=\frac{1}{N}$. Note that by assumption $\varepsilon\leq\frac{1}{100}$, we have $\delta\leq\varepsilon\leq\frac{1}{100}$.
@@
    Higher order derivative bounds.
@@
The following example shows that the local volume growth rate can still be positive in a small neighborhood of a repeller. Here the positivity of the local volume growth is a purely local phenomenon: although the curve is exponentially restricted to an increasingly small neighborhood of the repeller, the rapidly growing oscillation frequency produces a strong transverse amplification of derivatives. The resulting growth reflects a subtle balance between geometric restriction and oscillatory stretching, rather than uniform expansion.

### v2

@@ 4.2 Decomposition of the volume growth
We claim that $\mu$-almost every $x$ has a unique positive Lyapunov exponent (hence $\hat{\mu}$ is the unique unstable lift of $\mu$), due to the choice of $m_{n}$. Indeed, going by contradiction, if not, $\mu$ would have an ergodic component with two positive Lyapunov exponents and this ergodic component must support on the orbits of some periodic source $z$ with $\mu$ positive measure (see [29, Proposition 4.4] or [9, Fact in Claim 7.3]).
Let $r>0$ be small enough such that if $y\notin\cup_{i}B\left(f^{i}(z),r_{z}\right)$, then $f^{k}(y)\notin B(z,r)$ for all $k\in\mathbb{N}$. Indeed, it is enough to choose $r>0$ small enough such that for $0\leq j\leq T_{z}-1$ ($T_{z}$ denotes the period of $z$),
$f^{j}\left(B\left(z,r\right)\right)\subset B\left(f^{j}(z),r_{z}\right).$ (23)
To see this, we write $k=m\cdot T_{z}-j$ for some $0\leq j\leq T_{z}-1$. Since $y\notin B\left(f^{j}(z),r_{z}\right)$, by the choice of $r_{z}$ (see the second property in Definition 4.7),
$f^{j}\circ f^{k}(y)=f^{mT_{z}}(y)\notin B\left(f^{j}(z),r_{z}\right).$
By (23),
$f^{k}(y)\notin B\left(z,r\right).$
Since $\mu$ is not zero measure, we may assume for all large $n$,
$m_{n}+100<n.$
Since $x_{n,p}\in\sigma_{n}^{m_{n}}$, by the definition of $m_{n}$, there must exist some $0\leq i\leq m_{n}$ such that
$f^{i-1}(x_{n,p})\in B\left(f^{i-1}(z),r_{z}\right),\quad f^{i}(x_{n,p})\notin B\left(f^{i}(z),r_{z}\right).$
By the choice of $r_{z}$ (see the third property in Definition 4.7), we then have
$f^{i}(x_{n,p})\notin\bigcup_{i}B\left(f^{i}(z),r_{z}\right).$
@@
which is a contradiction. Hence $\mu$-almost every $x$ must have a unique positive Lyapunov exponent and $\hat{\mu}$ is the unique unstable lift of $\mu$.
@@
Now we use Proposition 4.10 to prove Theorem 2.1.
@@
Write
$\chi(f):=\max\left\{h_{\rm top}(f),\,\,\frac{\log\left\|Df\right\|}{r}\right\}+4\log\left(\log\left\|Df\right\|+\log\left\|Df^{-1}\right\|+2\right)+C_{r},\qquad\alpha\in[0,1]$
where $C_{r}$ is the universal constant depending only $r$ from Proposition 4.10.
@@
$\limsup_{n\to+\infty}\frac{1}{n}\log\sup_{\sigma}{\rm Vol}\left(f^{n}(\sigma)\right)\leq\chi(f).$ (24)
@@
$\limsup_{n\to+\infty}\frac{1}{n}\log\sup_{\sigma}{\rm Vol}\left(f^{n}(\sigma)\right)>\chi(f).$
@@
$\limsup_{k\to+\infty}\frac{1}{n_{k}}\log{\rm Vol}\left(f^{n_{k}}(\sigma_{n_{k}})\right)>\chi(f).$
Since each $\sigma_{n_{k}}$ can be decomposed into a uniformly bounded number of $\varepsilon_{f}$-bounded curves (see Lemma 6.7), we may, without loss of generality, assume that the sequence $\{\sigma_{n_{k}}\}$ itself consists of $\varepsilon_{f}$-bounded curves. Hence, by Proposition 4.10,
$\limsup_{k\to+\infty}\frac{1}{n_{k}}\log{\rm Vol}\left(f^{n_{k}}(\sigma_{n_{k}})\right)\leq\chi(f),$
@@
$\limsup_{n\to+\infty}\frac{1}{n}\log\sup_{\sigma}{\rm Vol}\left(f^{n}(\sigma)\right)\leq\max\left\{h_{\rm top}(f),\,\,\frac{\lambda^{+}(f)}{r}\right\}.$
@@ 5 Volume growth of the tangent cocycle
In this section, we study the relationship between the volume growth of the tangent cocycle and entropy. More precisely, we establish the chain of inequalities:
$h_{\rm top}(f)\leq\liminf_{n\to+\infty}\frac{1}{n}\log\int_{M}\left\|Df^{n}_{x}\right\|\,dx\leq\limsup_{n\to+\infty}\frac{1}{n}\log\int_{M}\left\|Df^{n}_{x}\right\|\,dx\leq\max\left\{h_{\rm top}(f),\,\,\frac{\lambda^{+}(f)}{r}\right\}.$
@@ 5.1 Upper bound of the volume growth rate
The following lemma asserts the existence of a point $z$ at which the exponential growth rate of $\int_{M}\left\|Df^{n}_{x}\right\|$ can be detected on arbitrarily small neighborhoods of $z$.
@@
Write
$\lambda(U):=\limsup_{n\to+\infty}\frac{1}{n}\log\int_{U}\left\|Df^{n}_{x}\right\|\,dx.$
@@
$\limsup_{n\to+\infty}\frac{1}{n}\log\left(\sum_{i=1}^{k}a_{n,i}\right)=\max_{1\leq i\leq k}\left(\limsup_{n\to+\infty}\frac{1}{n}\log a_{n,i}\right),$
@@
$\lambda(M)=\lambda\left(\bigcup_{i=1}^{k}U_{z_{i}}\right)=\max_{1\leq i\leq k}\lambda(U_{z_{i}})<\lambda(M)$
which is a contradiction. ∎
The main idea to prove Theorem 2.2 is to reduce the estimation of the volume growth of the tangent cocycle to that of one-dimensional submanifolds via Fubini’s theorem, which allows the two-dimensional integral of $\left\|Df^{n}_{x}\right\|$ to be decomposed into integrals along horizontal and vertical lines. The conclusion then follows by applying the curve volume growth estimate in Theorem 2.1.
@@
Let $\mathcal{A}:=\left\{\left(U_{i},\varphi_{i}\right):~\varphi_{i}:U_{i}\to[0,1]^{2}\right\}_{0\leq i<N}$ be a finite family of charts that cover $M$. Let $z$ be the point in Lemma 5.1. Without loss of generality, we assume $z\in U_{0}$ and $U_{0}$ is a ball at $z$ with sufficiently small diameter such that for any line $\gamma$ in $0,1]^{2}$, $\varphi^{-1}_{0}\left(\gamma\right)$ is contained in some $C^{r}$ embedded curve $\sigma:[0,1]\to M$ with $\left\|\sigma\right\|_{C^{r}}\leq 1$ and $\left\|d_{t}\sigma\right\|\geq 1/2,\,t\in[0,1]$. Let $K$ be a sufficiently large number (independent of $n$) such that
  • for any $0\leq i<N$ and any curve $\sigma$ on $U_{i}$,
@@
  • for any $0\leq i<N$ and any $n\geq 1$,
@@
    where $\left(x_{1},x_{2}\right)$ denotes the standard coordinates in $\mathbb{R}^{2}$ and
    $F^{n}:=\varphi_{i}\circ f^{n}\circ\varphi^{-1}_{0},\quad U_{0}^{i}:=\left\{x\in U_{0}:~f^{n}\left(x\right)\in U_{i}\right\}.$
Given $n\in\mathbb{N}$, we fix any $0\leq i<N$ such that
$\int_{U_{0}}\left\|Df^{n}_{x}\right\|\,d\leq N\cdot\int_{U_{0}^{i}}\left\|Df^{n}_{x}\right\|\,dx.$ (26)
@@
Applying Theorem 2.1 and recalling (25) and (26) , we have
$\limsup_{n\to+\infty}\frac{1}{n}\log\int_{M}\left\|Df^{n}_{x}\right\|\,dx$ $=\limsup_{n\to+\infty}\frac{1}{n}\log\int_{U_{0}}\left\|Df^{n}_{x}\right\|\,dx$
@@
$\leq\max\left\{h_{\rm top}(f),\,\,\frac{\lambda^{+}(f)}{r}\right\}.$
@@ 5.2 Lower bound of the volume growth rate
and whose cardinality is minimal among all finite subsets of $M$ with this property. Given a subset $\Omega$, recall (see [33] for background in ergodic theory) that a subset $S\subset\Omega$ is called a $(n,\varepsilon)$-separated set if for any $x,y\in S$ with $x\neq y$, there is $0\leq k<n$ such that $d(f^{k}(x),f^{k}(y))>\varepsilon$. Moreover, a $(n,\varepsilon)$-separated set $S$ is called maximal if its cardinality is larger than or equal to the cardinality of any other $(n,\varepsilon)$-separated set. Given a subset $\Omega$, let $S(n,\varepsilon,\Omega)$ denotes a maximal $(n,\varepsilon)$-separated set of $\Omega$. By maximality, the union of the dynamical balls $\{B(x,n,\varepsilon)\}_{x\in S(n,\varepsilon,\Omega)}$ covers $\Omega$.
@@
The following lemma, due to Przytycki, as an application of Pesin theory, provides a uniform lower bound on the integral of the derivative over dynamical balls centered at typical points w.r.t. a hyperbolic ergodic measure.
@@
[30, Lemma 2.1] Let $f$ be a $C^{r}$ ($r>1$) diffeomorphism on a compact surface $M$ and let $\mu$ be a hyperbolic ergodic measure. There are a subset $\Omega\subset M$ with $\mu\left(\Omega\right)>\frac{1}{2}$ and a constant $C>0$ such that for any $z\in\Omega$, any $n\in\mathbb{N}$ and any $0<\varepsilon<2026$,
@@
$\int_{M}\left\|Df^{n}_{x}\right\|\,dx$ $\geq\#S(n,2\varepsilon,\Omega)\cdot\min_{z\in S(n,2\varepsilon,\Omega)}\int_{B(z,n,\varepsilon)}\left\|Df^{n}_{x}\right\|\,dx$
@@
$\liminf_{n\to+\infty}\frac{1}{n}\log\int_{M}\left\|Df^{n}_{x}\right\|\,dx\geq\liminf_{n\to+\infty}\frac{1}{n}\log\#S(n,2\varepsilon,\Omega)-\log 2.$
Since $\mu(\Omega)>\frac{1}{2}$, we have $S(n,2\varepsilon,\Omega)\geq\#S_{\frac{1}{2}}(n,2\varepsilon)$. By Lemma 5.3 and the arbitrariness of $\varepsilon$, we then have
@@ 6 Appendix
Lemma 6.1.
Let $f$ be a homeomorphism on a compact metric space $(X,d)$. Given $n\in\mathbb{N}$ and $\varepsilon>0$, let $\left\{B(x_{i},n,\varepsilon)\right\}_{1\leq i\leq N}$ be a collection of mutually disjoint $(n,\varepsilon)$ dynamical balls. There is $0\leq k_{n}\leq n-1$ such that
$\sum_{i=1}^{N-1}d\left(f^{k_{n}}\left(x_{i}\right),f^{k_{n}}\left(x_{i+1}\right)\right)\geq\frac{\left(N-1\right)\varepsilon}{n}.$
Proof.
Since the dynamical balls $\left\{B(x_{i},n,\varepsilon)\right\}_{1\leq i\leq N}$ are mutually disjoint,
$\max_{0\leq k\leq n-1}d\left(f^{k}(x_{i}),f^{k}(x_{i+1})\right)\geq\varepsilon.$
Hence
$\sum_{i=1}^{N-1}\sum_{k=0}^{n-1}d\left(f^{k}(x_{i}),f^{k}(x_{i+1})\right)\geq\sum_{i=1}^{N-1}\max_{0\leq k\leq n-1}d\left(f^{k}(x_{i}),f^{k}(x_{i+1})\right)\geq(N-1)\varepsilon.$
Changing the order of summation yields:
$\sum_{k=0}^{n-1}\sum_{i=1}^{N-1}d\left(f^{k}(x_{i}),f^{k}(x_{i+1})\right)\geq(N-1)\varepsilon.$
Hence, there must exist at least one index $k_{n}\in\{0,1,\dots,n-1\}$ such that
$\sum_{i=1}^{N-1}d\left(f^{k_{n}}\left(x_{i}\right),f^{k_{n}}\left(x_{i+1}\right)\right)\geq\frac{\left(N-1\right)\varepsilon}{n}.$
Recall that $h_{\rm top}(f,\Omega)$ denotes the topological entropy on a subset $\Omega$, defined in the usual way by
$h_{\rm top}\left(f,\Omega\right):=\lim_{\varepsilon\to 0}\limsup_{n\to+\infty}\frac{1}{n}\log\#S(n,\varepsilon,\Omega)$
where $S(n,\varepsilon,\Omega)$ denotes a maximal $(n,\varepsilon)$-separated subset of $\Omega$.
Lemma 6.2.
Let $f$ be a $C^{1}$ diffeomorphism on a compact manifold $M$ and let $\sigma:[0,1]\to M$ be a $C^{1}$ curve. We have
$h_{\rm top}\left(f,\sigma\right)\leq\limsup_{n\to+\infty}\frac{1}{n}\log^{+}{\rm Vol}\left(f^{n}\left(\sigma\right)\right).$
Proof.
We may assume $h_{\rm top}\left(f,\sigma\right)>0$, otherwise the inequality is trivially true.
Given $0<\delta<h_{\rm top}\left(f,\sigma\right)/2$, let $\varepsilon>0$ be such that
$\limsup_{n\to+\infty}\frac{1}{n}\log\#S(n,2\varepsilon,\sigma)\geq h_{\rm top}\left(f,\sigma\right)-\delta>0$
where $S(n,2\varepsilon,\sigma)$ denotes a maximal $(n,2\varepsilon)$-separated set of $\sigma$. Write $S_{n}:=\#S(n,2\varepsilon,\sigma)$ for simplicity. We enumerate the points in $S(n,2\varepsilon,\sigma)$ as $\{x_{1},x_{2},\cdots x_{S_{n}}\}$ in such a way that $\sigma^{-1}\left(x_{i}\right)<\sigma^{-1}\left(x_{i+1}\right)$. Note that By definition, for any $z,z^{\prime}\in S(n,2\varepsilon,\sigma)$,
$B(z,n,\varepsilon)\cap B(z^{\prime},n,\varepsilon)=\emptyset.$
By Lemma 6.1, there is $0\leq k_{n}\leq n-1$ such that
$\sum_{i=1}^{S_{n}-1}d\left(f^{k_{n}}\left(x_{i}\right),f^{k_{n}}\left(x_{i+1}\right)\right)\geq\frac{\left(S_{n}-1\right)\varepsilon}{n}.$
As a consequence,
${\rm Vol}\left(f^{k_{n}}\left(\sigma\right)\right)\geq\frac{\left(S_{n}-1\right)\varepsilon}{n}.$
Hence, for all large $n$,
$\frac{1}{k_{n}}\log^{+}{\rm Vol}\left(f^{k_{n}}\left(\sigma\right)\right)\geq\frac{n}{k_{n}}\cdot\left(\frac{1}{n}\log\frac{\left(S_{n}-1\right)\varepsilon}{n}\right)\geq\frac{1}{n}\log\frac{\left(S_{n}-1\right)\varepsilon}{n}.$
Note that $k_{n}\to+\infty$ as $n\to+\infty$. Putting $\limsup_{n\to+\infty}$” on both sides, we then have
$\limsup_{n\to+\infty}\frac{1}{n}\log^{+}{\rm Vol}\left(f^{n}\left(\sigma\right)\right)\geq\limsup_{n\to+\infty}\frac{1}{k_{n}}\log^{+}{\rm Vol}\left(f^{k_{n}}\left(\sigma\right)\right)\geq h_{\rm top}\left(f,\sigma\right)-\delta.$
By the arbitrariness of $\delta$, we then have
$\limsup_{n\to+\infty}\frac{1}{n}\log^{+}{\rm Vol}\left(f^{n}\left(\sigma\right)\right)\geq h_{\rm top}\left(f,\sigma\right).$
The following lemma establishes an upper semi-continuity property for integrals of sub-additive potentials along converging sequences of invariant measures. This result is a special case of [12, Lemma 2.3].
@@
Let $X$ be a compact metric space and let $\mu$ be a Borel probability measure on $X$. Let $\{\varphi_{n}\}_{n\geq 1}$ be a sequence of measurable functions on $X$ such that $\sup_{n\geq 1}\left\|\varphi_{n}\right\|_{\infty}<+\infty$. Then
$\int_{X}\limsup_{n\to+\infty}\varphi_{n}\,d\mu\geq\limsup_{n\to+\infty}\int_{X}\varphi_{n}\,d\mu.$
The following lemma (see the book [28, Lemma 3.33]) gives an abstract comparison between the pointwise exponential growth rate and the exponential growth rate of the integral. We provide a proof for completeness.
Lemma 6.5.
Let $(X,\mathcal{A},\mu)$ be a probability space and $\varphi_{n}:X\to(0,+\infty)$ a sequence of integrable functions. Then
$\limsup_{n\to+\infty}\frac{1}{n}\log\varphi_{n}(x)\leq\limsup_{n\to+\infty}\frac{1}{n}\log\int_{X}\varphi_{n}\,d\mu,$
for $\mu$-a.e. $x\in X$.
Proof.
Write
$\lambda:=\limsup_{n\to+\infty}\frac{1}{n}\log\int_{X}\varphi_{n}\,d\mu.$
Given $\varepsilon>0$, we have
$\int_{X}\varphi_{n}\,d\mu\geq\int_{A_{n}}\varphi_{n}\,d\mu\geq\mu(A_{n})\cdot e^{n(\lambda+\varepsilon)}$
where
$A_{n}:=\left\{x\in X:~\varphi_{n}(x)\geq e^{n(\lambda+\varepsilon)}\right\}.$
Rearranging and taking the limit yields:
$\limsup_{n\to+\infty}\frac{1}{n}\log\mu(A_{n})\leq\limsup_{n\to+\infty}\frac{1}{n}\log\int_{X}\varphi_{n}\,d\mu-(\lambda+\varepsilon)=-\varepsilon<0.$
Hence
$\sum_{n=1}^{+\infty}\mu(A_{n})<+\infty.$
By the Borel-Cantelli Lemma, for $\mu$-a.e. $x$, there exists $N(x)$ such that for all $n\geq N(x)$,
$\varphi_{n}(x)<e^{n(\lambda+\varepsilon)},$
which proves the lemma by the arbitrariness of $\varepsilon$. ∎
Lemma 6.6.
Let $f$ be a homeomorphism on a compact metric space $X$. Let $\mu$ be an ergodic probability measure and let $A$ be a subset with $\mu(A)>0$. Let $\varphi:X\to\mathbb{R}$ be a measurable function. If for any $x\in A$, there exists $n_{x}\in\mathbb{N}$ such that $\varphi(f^{-n}(x))=0$ for all $n\geq n_{x}$, then $\varphi(x)=0$ for $\mu$-a.e. $x$.
Proof.
Define
$Z:=\left\{x\in X\mid\exists N\in\mathbb{N}\text{ s.t. }\forall n\geq N,\varphi(f^{-n}(x))=0\right\}.$
By definition, $Z$ is $f$-invariant, and $A\subset Z$. Since $\mu(A)>0$, ergodicity implies $\mu(Z)=1$.
If $\mu(\{\varphi\neq 0\})>0$, the Poincaré recurrence theorem applied to $f^{-1}$ implies that for $\mu$-a.e. $x\in\{\varphi\neq 0\}$, the backward orbit $f^{-n}(x)$ returns to $\{\varphi\neq 0\}$ infinitely often. This contradicts $\mu(Z)=1$. Thus, $\varphi=0$ $\mu$-a.e. ∎
@@
Take $N=\left\lceil\frac{1}{\varepsilon}\right\rceil$ and define the step size $\delta=\frac{1}{N}$. Note that by assumption $\varepsilon\leq\frac{1}{100}$, we have $\delta\leq\varepsilon\leq\frac{1}{100}$.
@@
    Higher order derivative bound.
@@
The following example shows that the local volume growth rate can still be positive in a small neighborhood of a repeller. Here the positivity of the local volume growth is a purely local phenomenon: although the curve is exponentially restricted to an increasingly small neighborhood of the repeller, the rapidly growing oscillation frequency produces a strong transverse amplification of derivatives. The resulting growth reflects a subtle balance between geometric restriction (uniform expansion) and oscillatory stretching.